A jet-ski driver wants to head to an island in the St. Lawrence River that is 5.0 km [W50.0S] away. If he is travelling at a speed of 40.0km/h and the St. Lawrence is flowing 6.0 km/h [E]
a) in what direction should he head the jet-ski?
b) How long will it take him to reach the island?
1
Expert's answer
2019-07-17T09:51:24-0400

a) he should sail in a westerly direction

b) the required time is

"t = \frac{5}{40-6} = 0.15 (h)"

 or  8.8 min